HUBBLE SPACE TELESCOPE DAILY REPORT #1710 PERIOD COVERED: 0000Z (UTC) 09/03/96 - 0000Z (UTC) 09/04/96 Daily Status Report as of 248/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 6188 (WF/PC-2 Cycle 5 Calibration: = Dark) The WF/PC-2 was used to make internal dark calibration images. = The observations were completed as scheduled, and no problems were = reported. 1.2 Completed WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS = Galaxy Sample-Part 1 High) The WF/PC-2 was used to make snapshot survey observations of the = ultraluminous IRAS galaxy QDOTDELLI. These data will be searched for = multiple cores, which imply a merger has occurred, and for starburst = regions. The observations were executed as scheduled, and no anomalies = were noted. 1.3 Completed GHRS/1 6070 (Interstellar Tin: Probing Galactic = S-Process Element Enrichment) The High Resolution Spectrograph (Side One) was used to make Echell= e A observations of star HD144217 in order to probe for interstellar tin = and to study the galactic S-process neutron capture of element = enrichment. The observations were completed as planned, and no problems = were reported. 1.4 Completed Three Sets of WF/PC-2 6253 (Pure Parallel Archive = Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with flu= x excesses at 3000 A0 and galactic sources with excesses at 1600 A0. The = observations were completed as planned, and no problems were reported. 1.5 Completed WF/PC-2 6905 (WF/PC-2 Cycle 6 Calibration: Internal = Monitor) The WF/PC-2 was used to make several internal calibrations = observations including biases and internal flats as part of a Cycle 6 = calibrations procedure. A variety of internal exposures are run weekly = to verify and monitor the health of the camera. The observations were = executed as scheduled, and no anomalies were noted. 1.6 Completed WF/PC-2/GHRS/1 6743 (HST Far UV Imaging and Spectra of = Jupiter's Aurora) The WF/PC-2 and the High Resolution Spectrograph (Side One) were = both used to make observations of the auroral zones on the planet = Jupiter. The observations were executed as scheduled, and no anomalies = were noted. 1.7 Completed WF/PC-2 6756 (Measuring Proper Motions of Galactic = Microlenses) The WF/PC-2 was used to make astrometric proper motion measurements= of STAR-1804 as part of a study of galactic microlensing. The = observations were completed as scheduled, and no problems were reported. 1.8 Completed GHRS/2 6455 (Metal Diffusion and Radiative Levitation in = Hot Helium Rich White Dwarfs) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the hot, helium-rich, white dwarf star = HD149499 in order to study its metal diffusion and radiative levitation. = The observations were executed as planned, and no anomalies were = reported. 1.9 Completed FOS/BL 6498 (Accretion Streams in Magnetic Variables: = FOS Eclipse of 4 Rosa) The Faint Object Spectrograph (Blue Side) was used to make Rapid = Mode spectrographic observations of the accretion streams in the magnetic = variable star (i.e., cataclysmic variable) RXJ1940-10. This is a binary = system, and it contains a white dwarf star, an accretion disk, a mass = stream, and hot spot where the mass stream hits the disk and a companion = star. The observations were completed as planned and no anomalies were = reported. 1.10 Completed GHRS/2 6501 (Probing the Equatorial Ejecta of Eta = Carinae) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the equatorial ejecta from the Eta-Carinae = system. The results will be used to interpret Galileo's magnetosphere = and auroral data. The combined data set will yield the relative = intensity and spatial distribution of auroral activity. The observations = were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 9 Successful : 9 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: Nothing to report 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH